Suppose we have events A and B so that P(A) = 0.15 and P(B) = 0.30.

(a) Assuming A and B are disjoint, calculate P(A or B).

(b) Assuming events A and B are independent, calculate P(A or B).
Answer #1

a) P(A or B) = P(A) + P(B) = 0.15+0.30 = 0.45

b) P(A or B) = P(A) + P(B) - P( A and B) = 0.15+0.30- (0.15*0.30) = 0.405